OpenStudy (jzjust):

Find the indicated power using De Moivre's Theorem. (Express your fully simplified answer in the form a + bi.) (2-2i)^7

OpenStudy (irishboy123):

thematically, as in trying to turn it into something based around angle and magnitude, it goes like this \((2-2i)^7\) \(\Rightarrow 2^7(1-i)^7\) \(\Rightarrow 2^7 \sqrt{2}^7(\frac{1}{\sqrt2}-\frac{i}{\sqrt2} ~ i)^7 \) etc so you've got \( \dfrac{\pi}{4}\) in there and a magnitude...

Directrix (directrix):

1. Write 2 - 2i in r*(cosθ + i*sinθ ) form. 2. Apply DeMoivre's Theorem. If you have difficulty, look at examples and discussion here: http://tinyurl.com/bg9pt9q

Directrix (directrix):

As a last resort, look at the worked problem here : http://tinyurl.com/hjsbfmd